OpenStudy (anonymous):

Identify the equation or coordinates of the following conic and its equation. Identify the center of the circle 34 = (x - 4)2 + (y + 7)2 Coordinates are ( , )

OpenStudy (valpey):

When a circle is centered at 0,0 its equation is of the form: \[r^2=x^2+y^2\] where r is the radius of the circle. When you displace x or y by x-a or x-b you just move the circle to a place centered at (a,b)

OpenStudy (anonymous):

can you give a quick example please, because I still don't quite understand

OpenStudy (anonymous):

you want \[34 = (x - 4)^2 + (y + 7)^2\] to look like \[r^2=(x-h)^2+(y-k)^2\] so you can read the answer right off the equation in \[r^2=(x-h)^2+(y-k)^2\] the radius is \(r\) and the center is \((h,k)\) write your equation as \[34=(x-4)^2+(y-(-7))^2\] and read the answer directly from the equation

OpenStudy (anonymous):

yes thx